# Quantum stochastic thermodynamics in the mesoscopic-leads formulation

### v1

@@ V.2.3 Probability to project into $m^{\mathrm{th}}$- eigenstate of $\rho(\tau)$
|$\ket{s_{m}}=\prod_{i=1}^{N_{\mathrm{S}}+L}(d_{i}^{\dagger})^{s_{m}^{i}}\ket{0}.$ (69)
In the following we are interested in computing the probability of the final state of the trajectory $\rho_{r}(\tau)$ to be projected into the eigenstate |$\ket{s^{o}_{m}}$ (here in the original/operational basis) of the ensemble average density matrix of the final state $\rho(\tau)$, with |$\ket{s^{o}_{m}}\bra{s^{o}_{m}}=\Pi_{m}^{\tau}$,
$p^{r\tau}_{m}=\mathrm{Tr}\left[\rho_{r}(\tau)\ket{s^{o}_{m}}\bra{s^{o}_{m}}\right].$ (70)
@@
$\begin{split}p^{r\tau}_{m}&=\mathrm{Tr}\left[\rho^{\prime}_{r}(\tau)\ket{s_{m}}\bra{s_{m}}\right].\end{split}$ (72)
@@
V.2.4 Probability for the backward trajectory to start in |$\ket{s_{m}}$
Now, what remains to be understood is the probability for the reverse trajectory to start in the eigenstate |$\ket{s_{m}}$ of the ensemble-averaged final state $\rho(\tau)$. This probability is simply given by the eigenvalue $p_{m}^{\tau}$ of $\rho(\tau)$ associated with the eigenstate |$\ket{s_{m}}$, defined in Eq. (68), and can thus be computed following the same method as described in Sec. V.2.1.
@@ V.3 Uncertainty and martingale entropy production
$\begin{split}e^{-\mathbf{c}^{\dagger}M_{1}\mathbf{c}}e^{-\mathbf{c}^{\dagger}M_{2}\mathbf{c}}&=e^{-\mathbf{c}^{\dagger}\chi\mathbf{c}}\\ \end{split}$ (81)
@@
$\begin{split}\mathrm{Tr}\left[\rho_{1}\rho_{2}\right]&=\frac{\mathrm{det}\left(1+e^{-M_{1}}e^{-M_{2}}\right)}{Z_{1}Z_{2}}.\\ \end{split}$ (84)
@@ V.4 Verification of the integral fluctuation theorems for entropy production
is coupled to a fermionic resevoir, represented in the mesoscopic-leads formalism, as shown in Fig. 4, in the steady-state. For simplicity, we consider the case in which the reservoir is described by a flat spectral density, given by
@@
where $\Gamma$ denotes the coupling strength between the quantum dot and the reservoir and $\omega_{\mathrm{max}}$ is a hard cut-off. We choose the set of lead mode energies $\{\epsilon_{k}\}$, the couplings between the quantum dot and the lead modes $\{\kappa_{k}\}$ and the damping rates of the residual reservoirs $\{\gamma_{k}\}$ as described in Sec. II. Here, we choose the energies $\{\epsilon_{k}\}$ via linear discretization.
@@ VI Heat dissipation fluctuations in finite-time information erasure
$-\Delta Q\geq T\log(2),$ (87)
where $-\Delta Q$ is the heat dissipated into the bath during the erasure process. The limit can be saturated only by a reversible and isothermal process, which requires infinite time. Recently, several studies have unveiled corrections to Landauer’s principle that appear in finite-time protocols Miller et al. (2020); Zhen et al. (2021); Van Vu and Saito (2022); Dago and Bellon (2022); Rolandi and Perarnau-Llobet (2022), but these studies have mostly been limited to the regime of weak system-bath coupling or slow driving.
Here, we exploit our method to study how varying the driving speed of the process impacts the dissipated heat fluctuations during information erasure with strong system-bath coupling. We consider a bit of information encoded in the occupation of a single fermionic mode (bit mode). The bit is erased by manipulating its time-dependent Hamiltonian, $H_{\mathrm{S}}(t)=\epsilon(t)f^{\dagger}f$, while in contact with a heat bath at temperature $T$ and chemical potential $\mu$. As above, we model this heat bath by a mesoscopic lead.
@@
where the so-called internal average particle and energy currents in the mesoscopic-leads formalism, exchanged between the system and the lead modes in lead $\alpha$, are generally defined as Lacerda et al. (2023b)
@@
We find that the mean dissipated heat approaches Landauer’s bound when the driving speed is reduced, as expected (see Fig. 7 a)). Further, the quantum fidelity to the target state, which for fermionic Gaussian states is given by Lacerda et al. (2023a); Banchi et al. (2014)
@@
Fig. 7 c) shows a histogram of the heat distribution obtained by trajectory sampling, which reveals distinctly non-Gaussian statistics even under slow-driving conditions. Previous work has shown that non-Gaussian heat statistics appear during slow information erasure in the weak-coupling regime, whenever quantum coherence is generated along the protocol Miller et al. (2020); Van Vu and Saito (2022). Here, coherence is expected due to the strong system-bath interaction; however, in contrast to the findings in Ref. Miller et al. (2020), we observe a negative skewness in the heat distribution. In addition, we observe rare events with extremely large heat flow into the system, despite the majority of trajectories yielding heat flow into the bath as demanded by Ineq. (87). It is important to note that the model of the bath employed here differs fundamentally with that of Ref. Miller et al. (2020) and thus agreement with those results is not to be expected. In particular, the dissipators and the Hamiltonian do not commute even though together they bring the system to equilibrium for large enough leads Lacerda et al. (2023a). Fig. 7 c) thus underlines the novel and nontrivial features of heat fluctuations that can emerge at strong system-bath coupling.
@@ VII Discussion
In this work, we have extended the continuous-time trajectory description of quantum stochastic thermodynamics to the strong coupling regime by exploiting a Markovian embedding. In particular, we have presented a comprehensive analysis of the conditional dynamics for the mesoscopic-leads approach, applied to non-interacting fermionic systems. We explicitly allow for imperfect detection and partial monitoring, enabling the exploration of various measurement configurations. Our analysis, based on the framework of quantum stochastic thermodynamics, yields thermodynamically consistent analytic expressions for particle and energy currents exchanged between the extended system and the residual reservoirs. We emphasize that these external currents correspond to those observable in experiments. They remain valid even in the strong coupling regime, where the differentiation between internal and external currents becomes pronounced.
@@ Appendix A Connection between $C$ and $n_{s}$
Given a fermionic Gaussian state $\rho$, of dimension $(2^{N}\times 2^{N})$, with eigenstates {|$\{\ket{s^{o}_{k}}\}$ and eigenvalues $\{\lambda_{s_{k}}\}$, its covariance matrix $C$, of dimension $(N\times N)$, is
$\begin{split}C_{ij}&=\mathrm{Tr}\left[\rho c_{j}^{\dagger}c_{i}\right]\\ &=\mathrm{Tr}\left[\sum_{k=1}^{2^{N}}\lambda_{s_{k}}\ket{s^{o}_{k}}\bra{s^{o}_{k}}c_{j}^{\dagger}c_{i}\right]\\ &=\sum_{k=1}^{2^{N}}\lambda_{s_{k}}\mathrm{Tr}\left[\ket{s^{o}_{k}}\bra{s^{o}_{k}}c_{j}^{\dagger}c_{i}\right]\\ &=\sum_{k=1}^{2^{N}}\lambda_{s_{k}}C^{s_{k}}_{ij}.\end{split}$ (96)
@@ C.1 Conditional dynamics
$\begin{split}\mathrm{d}\rho_{r}=&-i\left[H,\rho_{r}\right]\mathrm{d}t-\frac{1}{2}\sum_{k=1}^{L}\sum_{\sigma\in\{+,-\}}\mathcal{H}\left[{L^{\sigma}_{k}}^{\dagger}L^{\sigma}_{k}\right]\rho_{r}\mathrm{d}t\\ &+\sum_{k=1}^{L}\sum_{\sigma\in\{+,-\}}(1-\Lambda_{k}^{\sigma})\mathcal{G}\left[L^{\sigma}_{k}\right]\rho_{r}\mathrm{Tr}\left[{L^{\sigma}_{k}}^{\dagger}L^{\sigma}_{k}\rho_{r}\right]\mathrm{d}t\\ &+\sum_{k=1}^{L}\sum_{\sigma\in\{+,-\}}\Lambda_{k}^{\sigma}\mathcal{G}\left[L^{\sigma}_{k}\right]\rho_{r}\mathrm{d}N^{\sigma}_{k}.\end{split}$ (105)
It follows that the covariance matrix evolves, conditioned on the measurement record $r$, as
$\begin{split}\mathrm{d}C^{r}_{ij}(t)&=i\left\langle\left[H,c_{j}^{\dagger}c_{i}\right]\right\rangle_{r}\mathrm{d}t\\ &-\frac{1}{2}\sum_{k=1}^{L}\sum_{\sigma\in\{+,-\}}\left\langle\mathcal{H}^{+}\left[{L^{\sigma}_{k}}^{\dagger}L^{\sigma}_{k}\right]c^{\dagger}_{j}c_{i}\right\rangle_{r}\mathrm{d}t\\ &+\sum_{k=1}^{L}\sum_{\sigma\in\{+,-\}}(1-\Lambda_{k}^{\sigma})\left\langle\mathcal{G}^{+}\left[L^{\sigma}_{k}\right]c_{j}^{\dagger}c_{i}\right\rangle_{r}\langle{L^{\sigma}_{k}}^{\dagger}L^{\sigma}_{k}\rangle_{r}\mathrm{d}t\\ &+\sum_{k=1}^{L}\sum_{\sigma\in\{+,-\}}\Lambda_{k}^{\sigma}\left\langle\mathcal{G}^{+}\left[L^{\sigma}_{k}\right]c_{j}^{\dagger}c_{i}\right\rangle_{r}\mathrm{d}N_{k}^{\sigma}.\end{split}$ (106)
@@ C.1.1 No jump-conditioned evolution
$\begin{split}\frac{\mathrm{d}p}{\mathrm{d}t}&=-p\sum_{k=1}^{L}\sum_{\sigma\in\{+,-\}}\Lambda_{k}^{\sigma}\left\langle{L^{\sigma}_{k}}^{\dagger}{L^{\sigma}_{k}}\right\rangle_{r},\\ \end{split}$ (108)
@@ C.1.2 Jump-conditioned evolution
$\begin{split}\mathrm{d}C^{r}_{ij}(t_{J_{m}})=&\Lambda_{k_{m}}^{\sigma_{m}}\left\langle G^{+}\left[{L^{\sigma_{m}}_{k_{m}}}\right]c^{\dagger}_{j}c_{i}\right\rangle_{r}(t_{J_{m}}).\end{split}$ (111)
@@
$\mathrm{d}C^{r}(t_{J_{m}})=\bar{C}^{r}\left(\frac{\Lambda^{+}\mathrm{d}N_{+}(t_{J_{m}})}{\bar{C}^{r}}\right)\bar{C}^{r},$ (112)
@@
$\mathrm{d}C^{r}(t_{J_{m}})=-C^{r}\left(\frac{\Lambda^{-}\mathrm{d}N_{-}(t_{J_{m}})}{C^{r}}\right)C^{r}.$ (113)
@@
$\begin{split}\left(\frac{\Lambda^{+}\mathrm{d}N_{+}(t)}{\bar{C}^{r}}\right)_{kk}&=\frac{\Lambda^{+}_{k}\mathrm{d}N_{k+}(t)}{1-C^{r}_{kk}},\\ \left(\frac{\Lambda^{-}\mathrm{d}N_{-}(t)}{C^{r}}\right)_{kk}&=\frac{\Lambda^{-}_{k}\mathrm{d}N_{k-}(t)}{C^{r}_{kk}}.\end{split}$ (114)
@@ C.1.3 Stochastic ME for the covariance matrix
$\begin{split}\mathrm{d}C^{r}(t)=&\left[-(VC^{r}+C^{r}V^{\dagger})+C^{r}BC^{r}\right]\mathrm{d}t\\ &+\bar{C}^{r}F\left(1-\Lambda^{+}\right)\bar{C}^{r}\mathrm{d}t\\ &-C^{r}(\Gamma-F)\left(1-\Lambda^{-}\right)C^{r}\mathrm{d}t\\ &+\bar{C}^{r}\left(\frac{\Lambda^{+}\mathrm{d}N_{+}(t)}{\bar{C}^{r}}\right)\bar{C}^{r}\\ &-C^{r}\left(\frac{\Lambda^{-}\mathrm{d}N_{-}(t)}{C^{r}}\right)C^{r}.\end{split}$ (115)
@@ C.2 Particle current
$\begin{split}I^{r}_{N_{\alpha}}(t)\mathrm{dt}=&\mathrm{Tr}\left[C^{r}B_{\alpha}C^{r}-B_{\alpha}C^{r}\right]\mathrm{d}t\\ &+\mathrm{Tr}\left[\bar{C}^{r}\left(1-\Lambda^{+}_{\alpha}\right)F_{\alpha}\bar{C}^{r}\right]\mathrm{d}t\\ &-\mathrm{Tr}\left[C^{r}\left(1-\Lambda^{-}_{\alpha}\right)(\Gamma_{\alpha}-F_{\alpha})C^{r}\right]\mathrm{d}t\\ &+\mathrm{Tr}\left[\bar{C}^{r}\left(\frac{\Lambda^{+}\mathrm{d}N^{+}_{{\alpha}}(t)}{\bar{C}^{r}}\right)\bar{C}^{r}\right]\\ &-\mathrm{Tr}\left[C^{r}\left(\frac{\Lambda^{-}\mathrm{d}N^{-}_{{\alpha}}(t)}{C^{r}}\right)C^{r}\right],\end{split}$ (116)
@@
$\begin{split}\left(\Gamma_{\alpha}\right)_{kk},\left(F_{\alpha}\right)_{kk}=\begin{cases}\Gamma_{kk},F_{kk}&\text{if $k$-mode in lead $\alpha$,}\\ 0&\text{otherwise.}\\ \end{cases}\\ \left(1-\Lambda^{\sigma}_{\alpha}\right)_{kk}=\begin{cases}1-\Lambda^{\sigma}_{kk}&\text{if $k$-mode in lead $\alpha$,}\\ 0&\text{otherwise.}\\ \end{cases}\end{split}$ (117)
@@
$\begin{split}\left(\frac{\Lambda^{+}\mathrm{d}N_{+,\alpha}(t)}{\bar{C}^{r}}\right)_{kk}&=\begin{cases}\frac{\Lambda^{+}_{k}\mathrm{d}N_{k+,\alpha}(t)}{\bar{C}^{r}_{kk}}&\text{ if $k$-mode in lead $\alpha$,}\\ 0&\text{ otherwise.}\end{cases}\\ \left(\frac{\Lambda^{-}\mathrm{d}N_{-,\alpha}(t)}{C}\right)_{kk}&=\begin{cases}\frac{\Lambda^{-}_{k}\mathrm{d}N_{k-,\alpha}(t)}{C^{r}_{kk}}&\text{ if $k$-mode in lead $\alpha$,}\\ 0&\text{ otherwise.}\end{cases}\end{split}$ (118)

### v2

@@ V.2.3 Probability to project into $m^{\mathrm{th}}$- eigenstate of $\rho(\tau)$
[…]
$|].p^{r\tau}_{m}=\mathrm{Tr}\left[\rho_{r}(\tau)\ket{s^{o}_{m}}\bra{s^{o}_{m}}\right$]. (68)
@@
[…]
@@
[…]
@@ V.3 Uncertainty and martingale entropy production
[…]
@@
[…]
@@ V.4 Verification of the integral fluctuation theorems for entropy production
is coupled to a fermionic reservoir, represented in the mesoscopic-leads formalism, as shown in Fig. 4, in the steady-state. For simplicity, we consider the case in which the reservoir is described by a flat spectral density, given by
@@
where $\Gamma$ denotes the coupling strength between the quantum dot and the reservoir and $\omega_{\mathrm{max}}$ is a hard cut-off. We choose the set of lead mode energies $\{\epsilon_{k}\}$, the couplings between the quantum dot and the lead modes $\{\kappa_{k}\}$ and the damping rates of the residual reservoirs $\{\gamma_{k}\}$ as described in Sec. II. Here, we choose the energies $\{\epsilon_{k}\}$ via linear discretization. For details on the construction of the Liouvillian governing the dynamics of the the extended system given these parameters, we refer to Sec. II. Trajectories are recorded using the TPM scheme, with projective measurements taken in the eigenbasis of the steady-state of the extended system, both at the start and end of each trajectory, as described in Sec. V.
@@
We now briefly address why entropy production along trajectories remains nonzero and its average does not vanish, even in this example of an equilibrium process without external driving. This arises from the definition of total entropy production (Eq. (55)) under a local GKSL master equation, where bath-induced transitions do not mediate between system eigenstates. As a result, a given relative surprisal in measurement outcomes ($\log p_{n}^{0}/p_{m}^{\tau}$) cannot be uniquely linked to a specific heat change, unlike in a global GKSL framework. Consequently, even in an equilibrium state, there exists a distribution of total entropy flux. The nonzero average stems from the exclusion of measurement energy (Eq. (44)) in Eq. (55). However, including it would violate the integral fluctuation theorem in Eq. (53)Manzano and Zambrini (2022). The fact that the average measurement energy (Eq. (51)) is nonzero is a well-known issue in the mesoscopic leads framework and has been discussed before in Refs. Brenes et al. (2020); Lacerda et al. (2023b, 2024). However, this small discrepancy vanishes as the number of lead modes increases. We have included a more detailed discussion of this feature of the method in Appendix E.
@@ VI Heat dissipation fluctuations in finite-time information erasure
$\Delta Q\geq T\log(2),$ (85)
where $\Delta Q$ is the heat dissipated into the bath during the erasure process. The limit can be saturated only by a reversible and isothermal process, which requires infinite time. Recently, several studies have unveiled corrections to Landauer’s principle that appear in finite-time protocols Miller et al. (2020); Zhen et al. (2021); Van Vu and Saito (2022); Dago and Bellon (2022); Rolandi and Perarnau-Llobet (2022), but these studies have mostly been limited to the regime of weak system-bath coupling or slow driving.
Here, we exploit our method to study how varying the driving speed of the process impacts the dissipated heat fluctuations during information erasure with strong system-bath coupling. We consider a bit of information encoded in the occupation of a single fermionic mode (bit mode). The bit is erased by manipulating its time-dependent Hamiltonian, $H_{\mathrm{S}}(t)=\epsilon(t)f^{\dagger}f$, while in contact with a heat bath at temperature $T$ and chemical potential $\mu$. As above, we model this heat bath by a mesoscopic lead, as prescribed in Sec. II.
@@
where the so-called internal average particle and energy currents in the mesoscopic-leads formalism, exchanged between the system and the lead modes in lead $\alpha$, are generally defined as Lacerda et al. (2023a)
@@
We find that the mean dissipated heat approaches Landauer’s bound when the driving speed is reduced, as expected (see Fig. 7 a)). Further, the quantum fidelity to the target state, which for fermionic Gaussian states is given by Lacerda et al. (2024); Banchi et al. (2014)
@@
Fig. 7 c) shows a histogram of the heat distribution obtained by trajectory sampling, which reveals distinctly non-Gaussian statistics even under slow-driving conditions. Previous work has shown that non-Gaussian heat statistics appear during slow information erasure in the weak-coupling regime, whenever quantum coherence is generated along the protocol Miller et al. (2020); Van Vu and Saito (2022). Here, coherence is expected due to the strong system-bath interaction; however, in contrast to the findings in Ref. Miller et al. (2020), we observe a negative skewness in the heat distribution. In addition, we observe rare events with extremely large heat flow into the system, despite the majority of trajectories yielding heat flow into the bath as demanded by Ineq. (85). It is important to note that the model of the bath employed here differs fundamentally with that of Ref. Miller et al. (2020) and thus agreement with those results is not to be expected. In particular, the dissipators and the Hamiltonian do not commute even though together they bring the system to equilibrium for large enough leads Lacerda et al. (2024). Fig. 7 c) thus underlines the novel and nontrivial features of heat fluctuations that can emerge at strong system-bath coupling.
@@ VII Discussion
In this work, we have extended the continuous-time trajectory description of quantum stochastic thermodynamics to the strong coupling regime by exploiting a Markovian embedding. In particular, we have presented a comprehensive analysis of the conditional dynamics for the mesoscopic-leads approach, applied to non-interacting fermionic systems. We explicitly allow for imperfect detection and partial monitoring, enabling the exploration of various measurement configurations. Our analysis, based on the framework of quantum stochastic thermodynamics, yields thermodynamically consistent analytic expressions for particle and energy currents exchanged between the extended system and the residual reservoirs. We emphasize that these external currents correspond to those observable in experiments. They remain valid even in the strong coupling regime, where the differentiation between internal and external currents becomes pronounced. We emphasise that testing these predictions would not require an experimenter to identify or distinguish the lead from the residual reservoirs: these are simply theoretical constructs. Experimentally, one need only measure the integrated current with an ammeter and compare with the integrated current calculated using our method.
@@ Appendix A Connection between $C$ and $n_{s}$
[…]
@@ C.1 Conditional dynamics
$\begin{split}\mathrm{d}\rho_{r}=&-i\left[H,\rho_{r}\right]\mathrm{d}t-\frac{1}{2}\sum_{k=1}^{L}\sum_{\sigma\in\{+,-\}}\mathcal{H}\left[{L^{\sigma}_{k}}^{\dagger}L^{\sigma}_{k}\right]\rho_{r}\mathrm{d}t\\ &+\sum_{k=1}^{L}\sum_{\sigma\in\{+,-\}}(1-\Lambda_{k}^{\sigma})\mathcal{G}\left[L^{\sigma}_{k}\right]\rho_{r}\mathrm{Tr}\left[{L^{\sigma}_{k}}^{\dagger}L^{\sigma}_{k}\rho_{r}\right]\mathrm{d}t\\ &+\sum_{k=1}^{L}\sum_{\sigma\in\{+,-\}}\mathcal{G}\left[L^{\sigma}_{k}\right]\rho_{r}\mathrm{d}N^{\sigma}_{k},\end{split}$ (103)
where $\mathrm{E}[\mathrm{d}N^{\sigma}_{k}]=\Lambda_{k}^{\sigma}\mathrm{Tr}\left[{L^{\sigma}_{k}}^{\dagger}L^{\sigma}_{k}\rho_{r}\right]\mathrm{d}t$. It follows that the covariance matrix evolves, conditioned on the measurement record $r$, as
$\begin{split}\mathrm{d}C^{r}_{ij}(t)&=i\left\langle\left[H,c_{j}^{\dagger}c_{i}\right]\right\rangle_{r}\mathrm{d}t\\ &-\frac{1}{2}\sum_{k=1}^{L}\sum_{\sigma\in\{+,-\}}\left\langle\mathcal{H}^{+}\left[{L^{\sigma}_{k}}^{\dagger}L^{\sigma}_{k}\right]c^{\dagger}_{j}c_{i}\right\rangle_{r}\mathrm{d}t\\ &+\sum_{k=1}^{L}\sum_{\sigma\in\{+,-\}}(1-\Lambda_{k}^{\sigma})\left\langle\mathcal{G}^{+}\left[L^{\sigma}_{k}\right]c_{j}^{\dagger}c_{i}\right\rangle_{r}\langle{L^{\sigma}_{k}}^{\dagger}L^{\sigma}_{k}\rangle_{r}\mathrm{d}t\\ &+\sum_{k=1}^{L}\sum_{\sigma\in\{+,-\}}\left\langle\mathcal{G}^{+}\left[L^{\sigma}_{k}\right]c_{j}^{\dagger}c_{i}\right\rangle_{r}\mathrm{d}N_{k}^{\sigma}.\end{split}$ (104)
@@ C.1.1 No jump-conditioned evolution
[…]
@@ C.1.2 Jump-conditioned evolution
$\begin{split}\mathrm{d}C^{r}_{ij}(t_{J_{m}})=&\left\langle\mathcal{G}^{+}\left[{L^{\sigma_{m}}_{k_{m}}}\right]c^{\dagger}_{j}c_{i}\right\rangle_{r}(t_{J_{m}}).\end{split}$ (109)
@@
$\mathrm{d}C^{r}(t_{J_{m}})=\bar{C}^{r}\left(\frac{\mathrm{d}N_{+}(t_{J_{m}})}{\bar{C}^{r}}\right)\bar{C}^{r},$ (110)
@@
$\mathrm{d}C^{r}(t_{J_{m}})=-C^{r}\left(\frac{\mathrm{d}N^{-}(t_{J_{m}})}{C^{r}}\right)C^{r}.$ (111)
@@
$\begin{split}\left(\frac{\mathrm{d}N^{+}(t)}{\bar{C}^{r}}\right)_{kk}&=\frac{\mathrm{d}N^{+}_{k}(t)}{1-C^{r}_{kk}},\\ \left(\frac{\mathrm{d}N^{-}(t)}{C^{r}}\right)_{kk}&=\frac{\mathrm{d}N^{-}_{k}(t)}{C^{r}_{kk}}.\end{split}$ (112)
@@ C.1.3 Stochastic ME for the covariance matrix
$\begin{split}\mathrm{d}C^{r}(t)=&\left[-(VC^{r}+C^{r}V^{\dagger})+C^{r}BC^{r}\right]\mathrm{d}t\\ &+\bar{C}^{r}F\left(1-\Lambda^{+}\right)\bar{C}^{r}\mathrm{d}t\\ &-C^{r}(\Gamma-F)\left(1-\Lambda^{-}\right)C^{r}\mathrm{d}t\\ &+\bar{C}^{r}\left(\frac{\mathrm{d}N^{+}(t)}{\bar{C}^{r}}\right)\bar{C}^{r}\\ &-C^{r}\left(\frac{\mathrm{d}N^{-}(t)}{C^{r}}\right)C^{r}.\end{split}$ (113)
@@ C.2 Particle current
$\begin{split}I^{r}_{N_{\alpha}}(t)\mathrm{dt}=&\mathrm{Tr}\left[C^{r}B_{\alpha}C^{r}-B_{\alpha}C^{r}\right]\mathrm{d}t\\ &+\mathrm{Tr}\left[\bar{C}^{r}\left(1-\Lambda^{+}_{\alpha}\right)F_{\alpha}\bar{C}^{r}\right]\mathrm{d}t\\ &-\mathrm{Tr}\left[C^{r}\left(1-\Lambda^{-}_{\alpha}\right)(\Gamma_{\alpha}-F_{\alpha})C^{r}\right]\mathrm{d}t\\ &+\mathrm{Tr}\left[\bar{C}^{r}\left(\frac{\mathrm{d}N^{+}_{{\alpha}}(t)}{\bar{C}^{r}}\right)\bar{C}^{r}\right]\\ &-\mathrm{Tr}\left[C^{r}\left(\frac{\mathrm{d}N^{-}_{{\alpha}}(t)}{C^{r}}\right)C^{r}\right],\end{split}$ (114)
@@
[…]
@@
[…]